# Measuring Diversity in Synthetic Datasets

### v1

@@ Abstract
Large language models (LLMs) are widely adopted to generate synthetic datasets for various natural language processing (NLP) tasks, such as text classification and summarization. However, accurately measuring the diversity of these synthetic datasets—an aspect crucial for robust model performance—remains a significant challenge. In this paper, we introduce DCScore, a novel method for measuring synthetic dataset diversity from a classification perspective. Specifically, DCScore formulates diversity evaluation as a sample classification task, leveraging mutual relationships among samples. We further provide theoretical verification of the diversity-related axioms satisfied by DCScore, highlighting its role as a principled diversity evaluation method. Experimental results on synthetic datasets reveal that DCScore enjoys a stronger correlation with multiple diversity pseudo-truths of evaluated datasets, underscoring its effectiveness. Moreover, both empirical and theoretical evidence demonstrate that DCScore substantially reduces computational costs compared to existing approaches. Code is available at: https://github.com/BlueWhaleLab/DCScore.
@@ 1 Introduction
Large language models (LLMs) have shown exceptional performance across a range of fields, such as chatbots (Achiam et al., 2023), computer programming (Gu, 2023), and reasoning (Yuan et al., 2024). Inspired by their remarkable capacities, some research (Ye et al., 2022; Abdullin et al., 2024; Ding et al., 2024) employs LLMs as dataset generators to mitigate the shortage of training data. Although generated data facilitates model optimization, recent studies (Yu et al., 2024; Lee et al., 2023) suggest that a lack of diversity within the dataset—measured by the variation between samples (Long et al., 2024)—may lead to performance degradation in some scenarios. Although previous studies (Yu et al., 2024; Wang et al., 2022) leverage well-designed generation strategies to create highly diverse synthetic datasets, a crucial factor for measuring the diversity of these datasets has been overlooked. A principled diversity evaluation metric serves not only to guide LLM generators in creating more diverse data but also extends its utility to data selection (Cao et al., 2023), quantifying augmentation performance (Yang et al., 2024a), and assessing mode collapse (Dan Friedman & Dieng, 2023). Thus, a diversity evaluation method for synthetic datasets is becoming increasingly important.
Since the diversity evaluation of synthetic datasets remains under-explored, a natural solution is to directly employ diversity evaluation methods from the fields of natural language processing (NLP) (Khurana et al., 2023) and machine learning (ML) (Jordan & Mitchell, 2015). Specifically, efforts to measure diversity within these domains can be summarized into three categories: N-gram-based method (Zhu et al., 2018; Mishra et al., 2020), Reference-based method (Heusel et al., 2017; Cífka et al., 2018), and Transformation-based method (Du & Black, 2019; Zhang et al., 2024). The n-gram-based method assesses diversity through n-gram statistics, e.g., the distinct n-grams metric (Li et al., 2015) calculates the proportion of unique n-grams out of the total number of n-grams. This approach primarily focuses on the form differences of evaluated texts, often overlooking semantic aspects and offering limited flexibility for evaluators. To align the diversity criteria with human judgment, the reference-based method has emerged as a promising alternative. This approach employs a reference distribution or data as an approximation of human judgment and calculates the similarity between the evaluated data and the reference data (Holtzman et al., 2019). However, collecting reference data can be both time-consuming and may introduce potential biases.
@@
To sum up, the essence of diversity is associated with the identification of differences between samples, and the ability to distinguish these differences is a key element in the classification process (Quine, 1969). Motivated by this observation, we propose a synthetic dataset diversity evaluation method from a classification perspective, namely, DCScore. Notably, DCScore tackles the three aforementioned challenges. Firstly, DCScore treats the evaluation of each sample in the synthetic dataset as a distinct classification task, providing a holistic analysis. Secondly, theoretical verification demonstrates that DCScore satisfies four axioms outlined in (Leinster & Cobbold, 2012), including effective number, identical samples, symmetry, and monotonicity. Lastly, both empirical and theoretical evidence indicates that DCScore incurs lower computational costs compared to existing methods.
Our contributions can be summarized as follows:
@@
  • We theoretically validate that DCScore adheres to several intuitive axioms suggested by (Leinster & Cobbold, 2012), evidencing the superiority of DCScore.
  • Extensive experiments show that DCScore exhibits a stronger correlation with diversity parameters and human judgment compared to baseline metrics. We also perform a computational cost experiment to confirm the lower computational expense of DCScore.
@@ 2 Related Work
We give a brief literature review of diversity evaluation methods. Moreover, limited by the space, further related works on LLM dataset generators can be found in Appendix A.
@@ 2.1 Diversity Evaluation Methods
N-gram-based Methods. With the development of LLMs as dataset generators, the diversity evaluation of synthetic datasets has become a challenging task and remains under-explored in recent evaluation studies (Liang et al., 2022; tatsu lab, 2023). The most comparable diversity evaluation research can be traced back to studies in NLP and ML, which can be summarized into the n-gram-based method (Mishra et al., 2020), reference-based method (Heusel et al., 2017), and transformation-based method (Lai et al., 2020). The n-gram-based method is the most popular lexical diversity evaluation method, leveraging n-grams to capture differences in sentence form (Yu et al., 2024). Commonly used n-gram-based diversity metrics include distinct n-grams (distinct-n(Song et al., 2024), self-BLEU (Shu et al., 2019), and ROUGE-L (Wang et al., 2022; Padmakumar & He, 2023). However, this method has limitations, as it overlooks differences in other aspects such as semantics and style.
@@
Reference-based Methods. Diversity evaluation is inherently subjective, leading to a reliance on human judgment. Consequently, the reference-based method evaluates diversity by comparing the distribution of the evaluated data to that of a reference dataset (Heusel et al., 2017). MAUVE (Pillutla et al., 2021) exemplifies this idea by employing a divergence-based metric to capture correlations with human judgment. Regarding the natural language inference (NLI) training set as the reference dataset, (Stasaski & Hearst, 2022) first trains an NLI model to infer the relationship between pairs of generated texts and then calculates diversity based on these inference results. Due to the challenges in collecting reference datasets, a recent study (Le Bronnec et al., 2024) proposes evaluating diversity through precision and recall. Despite these advancements, the reference-based method remains significantly constrained because of the need for reference datasets.
@@
Transformation-based Methods. The transformation-based (Lee et al., 2023) method leverages well-designed models to generate representations of the evaluated data. Then, the diversity of these representations is summarized using techniques such as eigenvalue computation (Dan Friedman & Dieng, 2023) and clustering (Du & Black, 2019). Owing to the superior performance of representation learning, this method considers various aspects of the evaluated data, including semantics, form, and style, offering greater flexibility compared to the other two methods. However, its dependence on high-complexity summarization techniques, such as eigenvalue computation, limits its scalability for comprehensive evaluation tasks, such as evaluating the diversity of synthetic datasets.
@@
In summary, existing methods primarily focus on NLP and ML fields and are challenging to apply directly to overall dataset diversity evaluation. Different from the above-mentioned studies, our work is dedicated to the holistic diversity evaluation of synthetic datasets. Additionally, to ensure flexible evaluation, our work aims to evaluate diversity-sensitive components that impact the performance of trained models in terms of diversity.
3 Prelinimaries
@@ 3.2 Problem Formulation
where $\operatorname{Eval}$ is the diversity evaluation function, which takes $\{\mathcal{\tilde{T}}_{i}\}_{i=1}^{n}$ as input and outputs the diversity score $\operatorname{DiversityScore}$ of $\mathcal{D}$.
@@ 4 Present Work
Figure 1: Illustration of the computation of DCScore. DCScore consists of three stages: text representation, pairwise similarity, and diversity summarization.
@@ 4.1 DCScore: Measuring Diversity from a Classification Perspective
Due to the intrinsic nature of measuring sample differences in diversity evaluation, it is natural to evaluate diversity as a classification task. Consequently, we propose DCScore, which formulates diversity evaluation as a sample classification task. Specifically, the difference between samples can be measured through a $n$-classification task, where evaluating $n$ sample datasets involves $n$ $n$-classification tasks, with each sample corresponding to a distinct category. As shown in Figure 1, DCScore consists of three stages: text representation, pairwise similarity, and diversity summarization. According to the problem formulation in section 3.2, DCScore outputs the diversity of synthetic datasets by evaluating diversity-sensitive components.
Let $\mathcal{D}=\{\mathcal{T}_{i}\}_{i=1}^{n}=\{({x}_{i},{y}_{i})\}_{i=1}^{n}$ denote a synthetic dataset comprising $n$ input-output samples, and $\{\mathcal{\tilde{T}}_{i}\}_{i=1}^{n}$ represents the diversity-sensitive components. DCScore adheres to the paradigm of the transformation-based method to evaluate the diversity of $\mathcal{D}$. Specifically, given $\mathcal{\tilde{T}}_{i}$, DCScore first applies an embedding function $\Phi$ to extract the sample representation $\textbf{h}_{i}=\Phi(\mathcal{\tilde{T}}_{i})$. For all samples in $\mathcal{D}$, we obtain the sample representation matrix $\textbf{H}\in\mathbb{R}^{n\times d}$ across all samples, where $d$ denotes the dimension of sample representations. Subsequently, DCScore utilizes a kernel function $\operatorname{Kernel}$ to calculate a kernel matrix K, where $\textbf{K}\in\mathbb{R}^{n\times n}$ and entry $\textbf{K}[i,j]$ represents similarity between $\mathcal{\tilde{T}}_{i}$ and $\mathcal{\tilde{T}}_{j}$. From a classification perspective, $\textbf{K}[i,j]$ can be considered as the logit of $\mathcal{\tilde{T}}_{i}$ being classified into category $c_{j}$. where $c_{j}$ corresponds to $\mathcal{\tilde{T}}_{j}$. Formally, the aforementioned process can be formulated as follows:
@@
Based on the kernel matrix K, DCScore leverages a classification function with K, denoted as $f_{\textbf{K}}$, to compute the classification probability matrix $\textbf{P}\in\mathbb{R}^{n\times n}$. There are several potential choices for $f_{\textbf{K}}$, among which a natural option is the Softmax function. For $\mathcal{\tilde{T}}_{i}$, the probability that $\mathcal{\tilde{T}}_{i}$ is classified as category $c_{j}$ can be formulated as follows:
@@
According to (4), if the evaluated dataset exhibits high sample richness, indicating greater diversity, each sample is likely to be classified into its own category. Conversely, if diversity is low, all samples may be classified into a single category. Based on P, DCScore calculates diversity of $\mathcal{D}$ as the trace of P. The definition of diversity measurement from the classification perspective is as follows:
@@
Notably, the process described above is one way to implement of DCScore. Building on the core concept of DCScore, we have designed an alternative implementation, which is detailed in Appendix C.1.
@@ 4.2 Properties of DCScore
We provide theoretical proof that DCScore satisfies several axioms (Leinster & Cobbold, 2012) defined for a principled diversity metric. Specifically, DCScore meets four axioms: effective number, identical samples, symmetry, and monotonicity axioms. These guarantees ensure a reasonable and robust diversity evaluation. The matched axioms of our proposed method are outlined below, while their proofs are detailed in Appendix B due to space constraints.
@@
  • Symmetry: Diversity remains constant regardless of the order of the samples, exhibiting permutation invariance. Let $\pi$ denote the permutation function for the sample order, DCScore remains unchanged for any sample permutation of $\mathcal{D}$, i.e.,
@@
  • Monotonicity: The diversity of a dataset decreases as the similarity between its samples increases. Given two datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and a new sample $\mathcal{T}_{n+1}$, where the samples in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are entirely different, and $\operatorname{DCScore}(\mathcal{D}_{1})=\operatorname{DCScore}(\mathcal{D}_{2})=n$. If $\mathcal{T}_{n+1}$ is more similar to the samples in $\mathcal{D}_{2}$ than to those in $\mathcal{D}_{1}$ and is added to both datasets, then for the merged datasets $\mathcal{D}_{1}^{{}^{\prime}}$ and $\mathcal{D}_{2}^{{}^{\prime}}$, DCScore satisfies the following equation.
@@ 4.3 Complexity Analysis
Table 2: Comparison of Complexity analysis between DCScore and VendiScore. $\mathcal{O}_{kernel}$ represents the complexity of the kernel function.
@@
To better understand DCScore, we provide a brief analysis of its time complexity in Table 2. Since VendiScore is quite similar to DCScore, we compare the computational complexities between these two methods. Our analysis reveals that DCScore exhibits lower computation complexity under non-linear kernels compared to VendiScore.
Denoting $\mathcal{O}_{kernel}$ as the complexity associated with general kernels (i.e., kernels other than linear kernels), we analyze the complexity in the pairwise similarity and summarization stages. In the pairwise similarity stage, the computation of pairwise similarities results in a complexity of $\mathcal{O}(n^{2})$ for DCScore. When combined with the complexity $\mathcal{O}_{kernel}$ of the general kernel computation, DCScore exhibits a total complexity of $\mathcal{O}(n^{2}\cdot\mathcal{O}_{kernel})$ in this stage. In the summarization stage, DCScore has a complexity of $\mathbf{O}(n^{2})$ due to the softmax operation. Consequently, the overall complexity of DCScore for general kernels is $\mathcal{O}(n^{2}\cdot\mathcal{O}_{kernel}+n^{2})$. In contrast, VendiScore has a total complexity of $\mathcal{O}(n^{2}\cdot\mathcal{O}_{kernel}+n^{3})$, where the pairwise similarity stage is identical to that of DCScore, while the summarization stage incurs a complexity of $\mathcal{O}(n^{3})$ due to the eigenvalue computation. Thus, for general kernels, DCScore demonstrates lower complexity than VendiScore.
However, when the inner product is employed as the kernel function and $n\gg d$, VendiScore can significantly reduce the complexity by replacing the pairwise similarity $XX^{T}$ with $X^{T}X$, where $X\in\mathbb{R}^{n\times d}$. This results in complexities of $\mathcal{O}(d^{2}n)$ for the pairwise similarity stage and $\mathcal{O}(d^{3})$ for the summarization stage. In this regard, DCScore has a complexity of $\mathcal{O}(n^{2}d+n^{2})$, which is slightly worse than that of VendiScore. We can leverage efficient techniques, such as those proposed in (Shim et al., 2017), (Milakov & Gimelshein, 2018) and (Wen et al., 2023), to reduce the computational complexity of DCScore. Overall, compared to VendiScore, DCScore maintains lower complexity in most cases, as empirically validated in Section 5.3 and Appendix E.1. Although VendiScore has a lower complexity when the inner product is used as the kernel, experiments in the Appendix E.1 show that the computation times of VendiScore and DCScore are quite similar. In practical applications, different kernels keep computational complexity low, making them more useful than a single kernel.

### v2

@@ Abstract
Large language models (LLMs) are widely adopted to generate synthetic datasets for various natural language processing (NLP) tasks, such as text classification and summarization. However, accurately measuring the diversity of these synthetic datasets—an aspect crucial for robust model performance—remains a significant challenge. In this paper, we introduce DCScore, a novel method for measuring synthetic dataset diversity from a classification perspective. Specifically, DCScore formulates diversity evaluation as a sample classification task, leveraging mutual relationships among samples. We further provide theoretical verification of the diversity-related axioms satisfied by DCScore, highlighting its role as a principled diversity evaluation method. Experimental results on synthetic datasets reveal that DCScore enjoys a stronger correlation with multiple diversity pseudo-truths of evaluated datasets, underscoring its effectiveness. Moreover, both empirical and theoretical evidence demonstrate that DCScore substantially reduces computational costs compared to existing methods. Code is available at: https://github.com/bluewhalelab/dcscore.
@@ 1 Introduction
Large language models (LLMs) have shown exceptional performance across a range of fields, such as chatbots (Achiam et al., 2023), computer programming (Gu, 2023), and reasoning (Yuan et al., 2024). Inspired by their remarkable capacities, some research (Ye et al., 2022; Abdullin et al., 2024; Ding et al., 2024) employs LLMs as dataset generators to mitigate the shortage of training data. Although generated data facilitates model optimization, recent studies (Yu et al., 2024; Lee et al., 2023) suggest that a lack of diversity within the dataset—measured by the variation between samples (Long et al., 2024)—may lead to performance degradation in some scenarios. Although previous studies (Yu et al., 2024; Wang et al., 2022) leverage well-designed generation strategies to create highly diverse synthetic datasets, they consistently neglect to evaluate the diversity of these datasets. Additionally, a principled diversity evaluation metric serves not only to guide LLM generators in creating more diverse data but also extends its utility to data selection (Cao et al., 2023), quantifying augmentation performance (Yang et al., 2024a), and assessing mode collapse (Dan Friedman & Dieng, 2023). Thus, a diversity evaluation method for synthetic datasets is becoming increasingly important.
Since the diversity evaluation of synthetic datasets remains under-explored, a natural solution is to directly employ diversity evaluation methods from related fields, such as natural language processing (NLP) (Khurana et al., 2023) and machine learning (ML) (Jordan & Mitchell, 2015). Specifically, efforts to measure diversity within these domains can be summarized into three categories: N-gram-based method (Zhu et al., 2018; Mishra et al., 2020), Reference-based method (Heusel et al., 2017; Cífka et al., 2018), and Transformation-based method (Du & Black, 2019; Zhang et al., 2024). The n-gram-based method evaluates diversity through n-gram statistics, e.g., the distinct-n metric (Li et al., 2015) calculates the proportion of unique n-grams out of the total number of n-grams. This approach primarily focuses on the form differences of evaluated texts, often overlooking semantic aspects. To align the diversity criteria with human judgment, the reference-based method has emerged as a promising alternative. This approach employs a reference distribution or data as an approximation of human judgment and calculates the similarity between the evaluated data and the reference data (Holtzman et al., 2019). However, collecting reference data can be both time-consuming and may introduce potential biases.
@@
To sum up, the essence of diversity is associated with the identification of differences between samples, and the ability to distinguish these differences is a key element in the classification process (Quine, 1969). Motivated by this observation, we propose a synthetic dataset diversity evaluation method from a classification perspective, namely, DCScore. Notably, DCScore tackles the three aforementioned challenges. Firstly, DCScore treats the evaluation of each sample in the synthetic dataset as an independent classification task, providing a holistic analysis. Secondly, theoretical verification demonstrates that DCScore satisfies four axioms outlined in (Leinster & Cobbold, 2012), including effective number, identical samples, symmetry, and monotonicity. Lastly, both empirical and theoretical evidence suggest that DCScore effectively evaluates the diversity of synthetic datasets, while demonstrating lower computational costs compared to existing methods. Our contributions can be summarized as follows:
@@
  • We theoretically validate that DCScore adheres to several intuitive axioms suggested by (Leinster & Cobbold, 2012), demonstrating its superiority. Additionally, we theoretically validate the lower complexity of DCScore under general kernels.
  • Extensive experiments show that DCScore exhibits a stronger correlation with multiple diversity pseudo-truths compared to baseline metrics. We also perform a computational cost experiment to confirm the lower computational cost of DCScore.
@@ 2 Related Work
We give a brief literature review of diversity evaluation methods. Limited by space, further related works on LLM dataset generators and application of diversity evaluation methods can be found in Appendix A.
@@ 2.1 Diversity Evaluation Methods
With the development of LLMs as dataset generators, the diversity evaluation of synthetic datasets has become a challenging task and remains under-explored in recent evaluation studies (Liang et al., 2022; tatsu lab, 2023; Ribeiro et al., 2020). The most comparable diversity evaluation research can be traced back to studies in NLP and ML, which can be summarized into the n-gram-based method (Mishra et al., 2020), reference-based method (Heusel et al., 2017), and transformation-based method (Lai et al., 2020).
N-gram-based Methods. The n-gram-based method is the most popular lexical diversity evaluation method, leveraging n-grams to capture differences in sentence form (Yu et al., 2024). Commonly used n-gram-based diversity metrics include distinct n-grams (distinct-n(Song et al., 2024), self-BLEU (Shu et al., 2019), and ROUGE-L (Wang et al., 2022; Padmakumar & He, 2023). However, this type of method only captures differences in text form, thereby overlooking differences in other aspects such as semantics and style.
@@
Reference-based Methods. Diversity evaluation is a subjective task, leading to a reliance on human judgment. Consequently, the reference-based method evaluates diversity by comparing the distribution of the evaluated data to that of a reference dataset (Heusel et al., 2017). MAUVE (Pillutla et al., 2021) exemplifies this idea by employing a divergence-based metric to capture correlations with human judgment. Regarding the natural language inference (NLI) training set as the reference dataset, (Stasaski & Hearst, 2022) first trains an NLI model to infer the relationship between pairs of generated texts and then calculates diversity based on these inference results. Due to the challenges in collecting reference datasets, recent studies (Kynkäänniemi et al., 2019; Le Bronnec et al., 2024) propose evaluating diversity through precision and recall. (Naeem et al., 2020) introduces density and coverage as solutions to the susceptibility of precision and recall to outliers. Despite these advancements, the reference-based method remains significantly constrained because of the need for reference datasets.
@@
Transformation-based Methods. The transformation-based (Lee et al., 2023) method leverages well-designed models to generate representations of the evaluated data. Then, the diversity of these representations is summarized using techniques such as clustering (Du & Black, 2019) and eigenvalue computation (Dan Friedman & Dieng, 2023) of VendiScore (Dan Friedman & Dieng, 2023). In line with VendiScore (Dan Friedman & Dieng, 2023), RKE (Jalali et al., 2023) introduces an information-theoretic method for evaluating diversity in multimodal distributions, while FKEA (Ospanov et al., 2024) enhances the computational efficiency of both VendiScore and RKE. Owing to the superior performance of representation learning, this type of method considers various aspects of the evaluated data, including semantics, form, and style, offering greater flexibility compared to the other two methods. However, its dependence on high-complexity summarization techniques, such as eigenvalue computation, limits its scalability in evaluating the diversity of synthetic datasets.
@@
In summary, existing methods primarily focus on NLP and ML fields and are challenging to apply directly to synthetic dataset diversity evaluation. Different from the above-mentioned studies, our work is dedicated to the holistic diversity evaluation of synthetic datasets. Additionally, to ensure flexible evaluation, our work aims to evaluate diversity-sensitive components that impact the performance of trained models in terms of diversity.
3 Preliminaries
@@ 3.2 Problem Formulation
where $\operatorname{Eval}(\cdot)$ is the diversity evaluation function, which takes $\{\mathcal{\tilde{T}}_{i}\}_{i=1}^{n}$ as inputs and outputs the diversity score.
@@ 4 Present Work
Figure 1: The overview of DCScore. DCScore consists of text representation, pairwise similarity, and diversity summarization stages.
@@ 4.1 DCScore: Measuring Diversity from a Classification Perspective
Due to the intrinsic nature of measuring sample differences in diversity evaluation, it is natural to evaluate diversity as a classification task. Consequently, we propose DCScore, which formulates diversity evaluation as a sample classification task. Specifically, the difference between samples can be measured through an $n$-classification task, where evaluating $n$ sample datasets involves $n$ $n$-classification tasks, with each sample corresponding to a distinct category. As shown in Figure 1, DCScore consists of three stages: text representation, pairwise similarity, and diversity summarization. According to the problem formulation in section 3.2, DCScore outputs the diversity of synthetic datasets by evaluating diversity-sensitive components.
Let $\mathcal{D}=\{\mathcal{T}_{i}\}_{i=1}^{n}=\{({x}_{i},{y}_{i})\}_{i=1}^{n}$ denote a synthetic dataset comprising $n$ input-output samples, and $\{\mathcal{\tilde{T}}_{i}\}_{i=1}^{n}$ represents the diversity-sensitive components. DCScore adheres to the paradigm of the transformation-based method to evaluate the diversity of $\mathcal{D}$. Specifically, given $\mathcal{\tilde{T}}_{i}$, DCScore first applies an embedding function $\Phi$ to extract the sample representation $\textbf{h}_{i}=\Phi(\mathcal{\tilde{T}}_{i})$. For all samples in $\mathcal{D}$, we obtain the sample representation matrix $\textbf{H}\in\mathbb{R}^{n\times d}$ across all samples, where $d$ denotes the dimension of sample representations. Subsequently, DCScore utilizes a kernel function $\operatorname{Kernel}(\cdot)$ to calculate a kernel matrix K, where $\textbf{K}\in\mathbb{R}^{n\times n}$ and entry $\textbf{K}[i,j]$ represents similarity between $\mathcal{\tilde{T}}_{i}$ and $\mathcal{\tilde{T}}_{j}$. From a classification perspective, $\textbf{K}[i,j]$ can be considered as the logit of $\mathcal{\tilde{T}}_{i}$ being classified into category $c_{j}$. where $c_{j}$ corresponds to $\mathcal{\tilde{T}}_{j}$. Thus, a sample that is highly distinguishable from others (i.e., correctly classified into its ‘own’ category with high probability) contributes more to the overall diversity score. Formally, the aforementioned process can be formulated as follows:
@@
Based on K, DCScore leverages a classification function with K, denoted as $f_{\textbf{K}}(\cdot)$, to compute the classification probability matrix $\textbf{P}\in\mathbb{R}^{n\times n}$. Here, a natural option for $f_{\textbf{K}}(\cdot)$ is the Softmax function. For $\mathcal{\tilde{T}}_{i}$, the probability that $\mathcal{\tilde{T}}_{i}$ is classified as category $c_{j}$ can be formulated as follows:
@@
According to Eq. (4), if the evaluated dataset exhibits high sample richness, indicating greater diversity, each sample is likely to be classified into its own category. If the diversity is low, all samples will be randomly classified. Based on P, DCScore calculates diversity of $\mathcal{D}$ as the trace of P, which can be formulated as follows:
@@
Notably, the process described above is just one implementation of DCScore, with other potential implementations to be explored in future work.
@@ 4.2 Properties of DCScore
We provide theoretical proof that DCScore satisfies several axioms (Leinster & Cobbold, 2012) defined for a principled diversity metric. DCScore meets four axioms: effective number, identical samples, symmetry, and monotonicity axioms. These axioms ensure a reasonable and robust diversity evaluation. The matched axioms of DCScore are outlined below, while their proofs are detailed in Appendix B.
@@
  • Symmetry: Diversity remains constant regardless of the order of the samples, exhibiting permutation invariance. Let $\pi(\cdot)$ denote the permutation function for the sample order, DCScore remains unchanged for any sample permutation of $\mathcal{D}$, i.e.,
@@
  • Monotonicity: The diversity of a dataset decreases as the sample similarity increases. Given two datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, and a new sample $\mathcal{T}_{n+1}$, where the samples in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are entirely different, and $\operatorname{DCScore}(\mathcal{D}_{1})=\operatorname{DCScore}(\mathcal{D}_{2})=n$. If $\mathcal{T}_{n+1}$ is more similar to the samples in $\mathcal{D}_{2}$ than to those in $\mathcal{D}_{1}$ and is added to both datasets, then for the merged datasets $\mathcal{D}_{1}^{{}^{\prime}}$ and $\mathcal{D}_{2}^{{}^{\prime}}$, DCScore satisfies the following equation.
@@ 4.3 Complexity Analysis
We provide a time complexity analysis of DCScore in Table 2. We compare DCScore with VendiScore due to their similarity, finding that DCScore has lower computational complexity with non-linear kernels.
Denoting $\mathcal{O}_{kernel}$ as the complexity associated with general kernels (i.e., kernels other than linear kernels), we analyze the complexity in the pairwise similarity and summarization stages. In the pairwise similarity stage, the computation of pairwise similarities results in a complexity of $\mathcal{O}(n^{2})$ for DCScore. When combined with the complexity $\mathcal{O}_{kernel}$ of the general kernel computation, DCScore exhibits a total complexity of $\mathcal{O}(n^{2}\cdot\mathcal{O}_{kernel})$. In the summarization stage, DCScore has a complexity of $\mathcal{O}(n^{2})$ due to the softmax operation. Consequently, the overall complexity of DCScore for general kernels is $\mathcal{O}(n^{2}\cdot\mathcal{O}_{kernel}+n^{2})$. In contrast, VendiScore has a total complexity of $\mathcal{O}(n^{2}\cdot\mathcal{O}_{kernel}+n^{3})$, where the pairwise similarity stage is identical to that of DCScore, while the summarization stage incurs a complexity of $\mathcal{O}(n^{3})$ due to the eigenvalue computation. Thus, for general kernels, DCScore demonstrates lower complexity than VendiScore.
Table 2: Complexity analysis of DCScore and VendiScore. $\mathcal{O}_{kernel}$ represents the complexity of the kernel function.
@@
However, when the inner product is employed as the kernel function and $n\gg d$, VendiScore can significantly reduce the complexity by replacing the pairwise similarity $XX^{T}$ with $X^{T}X$, where $X\in\mathbb{R}^{n\times d}$. This results in complexities of $\mathcal{O}(d^{2}n)$ for the pairwise similarity stage and $\mathcal{O}(d^{3})$ for the summarization stage. In this regard, DCScore has a complexity of $\mathcal{O}(n^{2}d+n^{2})$, which is slightly worse than that of VendiScore. We can leverage efficient techniques, such as those proposed in (Shim et al., 2017), (Milakov & Gimelshein, 2018), and (Wen et al., 2023), to reduce the computational cost of DCScore. Compared to VendiScore, DCScore maintains lower complexity in most cases, as empirically validated in Section 5.3 and Appendix D.1. While VendiScore has lower complexity with the inner product kernel, experiments in Appendix D.1 indicate that its computation time is similar to DCScore. Ensuring low computational complexity across multiple kernels is more advantageous than achieving it with a single kernel.